Systems of Linear Equations word problem:
Can someone tell me how I set up this word problem please?
A boat traveled 27.5 mph with the current and 12.5 mph against the current. Find the speed of the boat in (still) water and the speed of the current?

Answer by richwmiller(17219)

b+c=27.5
b-c=12.5
2b=40
b=20 mph in still water
27.5-20=7.5 mph=current